OpenStudy (anonymous):

Solve x – 2y = 6 for y i got y=(6-x)/2

OpenStudy (anonymous):

i mean -2

OpenStudy (kinggeorge):

That looks correct to me. (with the -2)

jimthompson5910 (jim_thompson5910):

x - 2y = 6 -2y = 6 - x -2y = -x + 6 y = -x/(-2) + 6/(-2) y = x/2 - 3 y = (1/2)x - 3

OpenStudy (kinggeorge):

You could also try \[(x-6) \over 2\]By factoring out a \[{-1 \over -1}=1\]

OpenStudy (kinggeorge):

\[{6-x \over -2}={-1 \over -1}\;\;{x-6 \over 2}={x-6 \over 2}\]